OpenStudy (anonymous):

c(c+3)-c(c-4)=9c-16 simplify

OpenStudy (anonymous):

c=8 Just distribute then add like terms. It is really not difficult if you try

OpenStudy (anonymous):

i have a 27% in algebra and thats with me trying im re taking algebra next year and plus some of them i have answered already and am just seeing if people get same answer as me or what lol

OpenStudy (anonymous):

well do you want me to walk you throuhg the steps or no

OpenStudy (anonymous):

if you want you can i have one more practice problem that teacher wants me to do

OpenStudy (anonymous):

c^2 + 3c - c^2 + 4c = 9c - 16 distribute 7c = 9c - 16 add like terms -2c = -16 -9c from both sides c = 8 divide both sides by -2

OpenStudy (anonymous):

well thanks blexting she just did it for you just pay attention to what she did its the similar procedure for all of your problems

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

thanks